OpenStudy (anonymous):

evaluate the function y=-3(1/4)^x for x =-3 a. -3/16 b. 1/27 c. -192 d. 192

OpenStudy (anonymous):

y=-3((1/4)^(-3))=(-3)(4^3)

OpenStudy (anonymous):

(-3)(4^3)=-192